OpenStudy (hellokitty17):

The following triangle FOX is transformed using a rotation: Which could be coordinates of F'O'X' after the rotation? F=(1,1), O=(3,2), X=(0,3) F = (1, 1), O = (−3, −2), X = (0, −3) F=(1, −1), O=(3,−2), X=(0,−3) F=(−1,1), O=(−3, 2), X=(0,3)

OpenStudy (hellokitty17):

@izzy100200

OpenStudy (anonymous):

whats the answer to this first F=(1, −1), O=(3,−2), X=(0,−3)

OpenStudy (anonymous):

Who do you think deserves to win the war: the Greeks or the Trojans? Why?

OpenStudy (hellokitty17):

is that the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Who do you think deserves to win the war: the Greeks or the Trojans? Why?

OpenStudy (anonymous):

@HelloKitty17 come on i helped you